OpenStudy (anonymous):

sole this one if you canfind the equation of the line through (2,-3,5) and parallel to 5i-j+2k

OpenStudy (mertsj):

\[x=x _{0}+at \] \[y=y _{0}+at\] \[z=z _{0}+at\] \[r _{0}=2i-3j+5k\] \[v=5i-j+2k\] \[r=(2i-3j+5k)+t(5i-j+2k)\] \[r=(2+5t)i+(-3-t)j+(5+2t)k\]

OpenStudy (mertsj):

Parametic equations are: \[x=2+5t\] \[y=-3-t\] \[z=5+2t\]

OpenStudy (mertsj):

\[\frac{x-2}{5}=\frac{y+3}{-1}=\frac{z-5}{2}\]